# Deep learning for dynamic graphs: models and benchmarks

### v1

@@ IV-B2 Stacked architectures
$\begin{split}\tilde{\mathbf{H}}^{\ell}_{t}&=\text{{dgn}}^{\ell}(\tilde{\mathbf{H}}_{t}^{\ell-1})\\ \mathbf{H}_{t}^{\ell}&=\text{{update}}(\tilde{\mathbf{H}}^{\ell}_{t},\mathbf{H}_{t}^{\ell-1}).\end{split}$ (14)
where update is the sequence encoder.
Contrarily from previous works, Cini et al. [47] propose to first embed the history of the node time series into latent representations that encode the temporal dynamic of the system. Such representations are then processed leveraging multiple powers of a graph shift operator (e.g., graph laplacian or adjacency matrix) to encode the spatial dynamic of the system. Specifically, the authors propose to encode the temporal dynamic by means of an Echo State Networks (ESNs) [48, 49], a randomized recurrent neural networks, to efficiently compute node embedding and improve the scalability of DGNs for D-TDGs.
With the same aim of speeding up the dynamic graph processing, Micheli and Tortorella [50] propose DynGESN, an extension of the Graph Echo State Network [51] to the temporal domain. Specifically, DynGESN updates the embedding for a node $u$ at time $t$ as
$\mathbf{h}_{t}^{u}=(1-\gamma)\mathbf{h}_{t-1}^{u}+\gamma\text{tanh}\left(\mathbf{\Theta}_{i}\mathbf{h}^{u}_{t}+\sum_{v\in\mathcal{N}^{t}_{u}}\mathbf{\Theta}_{r}\mathbf{x}^{v}_{t-1}\right),$ (15)
@@ IV-B3 Meta architectures
$\begin{split}\mathbf{\Theta}_{t}&=\mathrm{GRU}(\mathbf{X}_{t},\mathbf{\Theta}_{t-1})\\ \mathbf{H}_{t}&=\mathrm{GCN}(\mathbf{X}_{t},\mathcal{E}_{t},\mathbf{\Theta}_{t})\end{split}$ (16)
@@ IV-B4 Autoencoder architectures
$\begin{split}\mathrm{encoder}:\,\mathbf{h}^{enc}_{t}&=\mathrm{LSTM}_{enc}(emb(\mathcal{G}_{t}),\mathbf{h}^{enc}_{t-1})\\ \mathrm{decoder}:\,\mathbf{h}^{dec}_{t}&=\mathrm{LSTM}_{dec}(\bar{\mathbf{A}}_{t-1},\mathbf{h}^{dec}_{t-1})\end{split}$ (17)
@@
$\mathbf{h}^{dec}_{t}=\mathrm{LSTM}_{dec}([\mathbf{h}_{t}^{*}||\bar{\mathbf{A}}_{t-1}],\mathbf{h}^{dec}_{t-1})$ (18)
where $\mathbf{h}_{t}^{*}=\sum_{i=t-w}^{t-1}\bar{\alpha}_{t}^{i}\mathbf{h}_{i}^{enc}$ is the attention distribution, the attention weights $\bar{\alpha}_{t}^{i}=\mathrm{softmax}(f(\mathbf{h}^{dec}_{t-1},\mathbf{h}^{enc}_{i}))$, and $f$ is a function, e.g., dot product or MLP.
@@
$size(\ell+1)\geq\rho\cdot size(\ell)$ (19)
@@ V-1 Integrated architectures
$\hat{\mathbf{x}}_{u}(t)=(1+\mathbf{w})\odot\mathbf{x}_{u}(t^{-}_{u})$ (20)
@@
Trivedi et al. [11] proposed DyRep, a framework that update the representation of a node as it appear in an event in the C-TDG. DyRep captures the continuous-time dynamics leveraging a temporal point process approach. A temporal point process is characterized by the conditional intensity function that models the likelihood of an event to happen given the previous events. DyRep considers only two event types: topological evolution (i.e., node/edge creation/deletion), and node interaction (i.e., related to activity between two nodes in the graphs). DyRep’s conditional intensity function, computed for an event between nodes $u$ and $v$ at time $t$, is:
$\lambda_{uv}^{k}(t)=f_{k}(g^{k}_{uv}(t^{-}))$ (21)
@@
$f_{k}(z)=\psi_{k}\log\left(1+\exp\left(\frac{z}{\psi_{k}}\right)\right)$ (22)
@@
$g^{k}_{uv}(t^{-})=\omega_{k}^{T}\cdot[\mathbf{x}_{u}(t^{-})||\mathbf{x}_{v}(t^{-})]$ (23)
@@
$\mathbf{h}_{u}(t)=\sigma(\mathbf{\Theta}_{i}\mathbf{h}_{u}^{loc}(t^{-})+\mathbf{\Theta}_{r}\mathbf{h}_{u}(t^{-}_{u})+\mathbf{\Theta}_{e}(t-t^{-}_{u}))$ (24)
where $h_{u}^{loc}(t^{-})\in\mathbb{R}^{d_{h}}$ is the representation of the aggregation of $u$’s direct neighbors, $t^{-}_{u}$ is the previous event involving node $u$, and $\mathbf{\Theta}_{i},\mathbf{\Theta}_{r}\in\mathbb{R}^{d_{h}\times d_{h}}$ and $\mathbf{\Theta}_{e}\in\mathbb{R}^{d_{h}}$ are learnable parameters. In Equation 24 the first addend propagates neighborhood information, the second self-information, while the third considers the exogenous force that may smoothly update node features during the interval time. To learn $\mathbf{h}_{u}^{loc}(t^{-})$, DyRep uses an attention mechanism similar to the one proposed in the GAT model by Veličković et al. [20]. In this case, the attention coefficient is parametrized by $\mathcal{S}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$, which is a stochastic matrix denoting the likelihood of communication between each pair of nodes. $\mathcal{S}$ is updated according to the conditional intensity function. The aggregated neighborhood representation is
$\mathbf{h}_{u}^{loc}(t^{-})=\max(\{\sigma(\alpha_{uv}(t)\cdot\mathbf{h}_{v}(t^{-}))\,|\,v\in\mathcal{N}_{u}^{t})\}),$ (25)
@@ V-2 Stacked architectures
$\mathbf{h}^{\ell}_{u}(t)=\text{{mlp}}^{\ell}_{2}(\mathrm{ReLU}(\text{{mlp}}^{\ell}_{1}([\hat{\mathbf{h}}(t)||\mathbf{x}_{u}])))$ (26)
where $\hat{\mathbf{h}}(t)$ is the hidden neighborhood representation obtained as
$\begin{split}\mathbf{q}(t)&=\left[\mathbf{Z}(t)\right]_{0}\mathbf{\Theta}_{q}\\ \mathbf{K}(t)&=\left[\mathbf{Z}(t)\right]_{1:n}\mathbf{\Theta}_{K}\\ \mathbf{V}(t)&=\left[\mathbf{Z}(t)\right]_{1:n}\mathbf{\Theta}_{V}\\ \hat{\mathbf{h}}(t)&=\mathrm{attn}(\mathbf{q}(t),\mathbf{K}(t),\mathbf{V}(t))\end{split}$ (27)
where $\mathbf{Z}(t)=\left[\mathbf{x}^{\ell-1}_{u}(t)||\Phi_{d}(0),...,\mathbf{x}^{\ell-1}_{v}(t)||\Phi_{d}(t-t_{v})],...\right]$ is the temporal feature matrix, with $v\in\mathcal{N}_{u}^{t}$ and $n$ the size of $u$’s neighborhood; $\mathbf{q}(t)$, $\mathbf{K}(t)$, and $\mathbf{V}(t)$ are the query, key and value projections of the matrix; and $\mathrm{attn}$ is an attention mechanism similar to GAT. The dimensional functional mapping $\Phi_{d}:t\rightarrow\mathbb{R}^{d_{h}}$ is defined as
$\Phi_{d}(t)=[\cos(\omega_{1}t)\sin(\omega_{1}t),...,\cos(\omega_{d}t)\sin(\omega_{d}t)]$ (28)
@@
$\begin{split}m_{u}(t)&=\mathrm{msg}_{src}(\mathbf{s}_{u}(t^{-}),\mathbf{s}_{v}(t^{-}),\Delta t,\mathbf{e}_{uv}(t))\\ m_{v}(t)&=\mathrm{msg}_{dst}(\mathbf{s}_{v}(t^{-}),\mathbf{s}_{u}(t^{-}),\Delta t,\mathbf{e}_{uv}(t)),\end{split}$ (29)
@@
$\mathbf{s}_{u}(t)=\mathrm{mem}(\bar{m}_{u}(t),\mathbf{s}_{u}(t^{-}))$ (30)
where $\bar{m}_{u}(t)$ represents the aggregated messages, and $\mathrm{mem}$ is an RNN. Lastly, the embedding module generates the representation for a node $u$ at time $t$ by exploiting the information stored in the memory module of the node itself and its neighborhood up to time $t$
$\mathbf{h}_{u}(t)=\sum_{v\in\mathcal{N}_{u}^{t}}f(\mathbf{s}_{u}(t),\mathbf{s}_{v}(t),\mathbf{x}_{u}(t),\mathbf{x}_{v}(t),\mathbf{e}_{uv})$ (31)
@@ V-3 Random walk based architectures
$\mathbb{P}(v)=\frac{\exp(\mathcal{T}(v)-\mathcal{T}(u))}{\sum_{v^{\prime}\in\mathcal{N}^{t}_{u}}\exp(\mathcal{T}(v^{\prime})-\mathcal{T}(u))},$ (32)
@@
$\mathbb{P}(v)=\frac{\delta(v,\mathcal{T}(v))}{\sum_{v^{\prime}\in\mathcal{N}^{t}_{u}}\delta(v^{\prime},\mathcal{T}(v^{\prime}))},$ (33)
where $\delta:\mathcal{V}\times\mathbb{R}^{+}\rightarrow\mathbb{Z}^{+}$ is a function that sorts temporal neighbors in descending order.
@@ V-4 Hybrid architectures
$\begin{split}\hat{\mathbf{h}}^{\ell}_{u}(t)&=\sum_{v\in\mathcal{N}_{u}^{t}}\text{{mlp}}^{\ell}_{agg}(\mathbf{h}^{\ell-1}_{v}(t)||\mathbf{e}_{uv})\alpha^{-\beta(t-t^{-})}\\ \mathbf{h}^{\ell}_{u}(t)&=\text{{mlp}}^{\ell}_{upd}(\mathbf{h}^{\ell-1}_{u}(t)||\hat{\mathbf{h}}^{\ell}_{u}(t))\end{split}$ (35)
@@ VI The benchmark problem
With this in mind, we designed three benchmarks to assess models that deal with spatio-temopral graphs, general D-TDGs, and C-TDGs. To do so, we extended the library PyDGN [^3] https://github.com/diningphil/PyDGN to the D-TDG learning setting to foster reproducibility and robustness of results. With the same aim, we developed a Pytorch Geometric [62] based framework to allow reproducible results in the continuous scenario. Lastly, in Table I we provide the community with a selection of datasets useful for benchmarking future works. An interest reader is referred to SNAP [63], TSL [64], and Network Repository [65] for a broader data collections.

### v2

@@ IV-B2 Stacked architectures
$\small\begin{split}\tilde{\mathbf{H}}^{\ell}_{t}&=\text{{dgn}}^{\ell}(\tilde{\mathbf{H}}_{t}^{\ell-1})\\ \mathbf{H}_{t}^{\ell}&=\text{{update}}(\tilde{\mathbf{H}}^{\ell}_{t},\mathbf{H}_{t}^{\ell-1}).\end{split}$ (14)
where update is the sequence encoder. Similarly Deng et al. [52] employed GCN to learn spatial dependencies and MLP as update function.
Contrarily from previous works, Cini et al. [53] propose to first embed the history of the node time series into latent representations that encode the temporal dynamic of the system. Such representations are then processed leveraging multiple powers of a graph shift operator (e.g., graph laplacian or adjacency matrix) to encode the spatial dynamic of the system. Specifically, the authors propose to encode the temporal dynamic by means of an Echo State Networks (ESNs) [54, 55], a randomized recurrent neural networks [^3] In randomized neural networks the hidden weights are randomly initialized and kept fixed after initialization. Only the weights in the final readout layer are learned, typically employing highly efficient methods like least-squared minimization [56]., to efficiently compute node embedding and improve the scalability of DGNs for D-TDG.
@@ IV-B3 Meta architectures
$\small\begin{split}\mathbf{\Theta}_{t}&=\mathrm{GRU}(\mathbf{X}_{t},\mathbf{\Theta}_{t-1})\\ \mathbf{H}_{t}&=\mathrm{GCN}(\mathbf{X}_{t},\mathcal{E}_{t},\mathbf{\Theta}_{t})\end{split}$ (15)
@@ IV-B4 Autoencoder architectures
$\small\begin{split}\mathrm{encoder}:\,\mathbf{h}^{enc}_{t}&=\mathrm{LSTM}_{enc}(emb(\mathcal{G}_{t}),\mathbf{h}^{enc}_{t-1})\\ \mathrm{decoder}:\,\mathbf{h}^{dec}_{t}&=\mathrm{LSTM}_{dec}(\bar{\mathbf{A}}_{t-1},\mathbf{h}^{dec}_{t-1})\end{split}$ (16)
@@
$\small\mathbf{h}^{dec}_{t}=\mathrm{LSTM}_{dec}([\mathbf{h}_{t}^{*}||\bar{\mathbf{A}}_{t-1}],\mathbf{h}^{dec}_{t-1})$ (17)
where $\mathbf{h}_{t}^{*}=\sum_{i=t-w}^{t-1}\bar{\alpha}_{t}^{i}\mathbf{h}_{i}^{enc}$ is the attention distribution, the attention weights $\bar{\alpha}_{t}^{i}=\mathrm{softmax}(f(\mathbf{h}^{dec}_{t-1},\mathbf{h}^{enc}_{i}))$, and $f$ is a function, e.g., dot product or MLP. Dyngraph2vec [60] introduces an analogous encoder-decoder approach featuring a deep architecture comprising dense and recurrent layers. This design facilitates the utilization of a more extended temporal evolution for predictions.
@@
$\small size(\ell+1)\geq\rho\cdot size(\ell)$ (18)
@@ V-1 Integrated architectures
$\small\hat{\mathbf{x}}_{u}(t)=(1+\mathbf{w})\odot\mathbf{x}_{u}(t^{-}_{u})$ (19)
@@
Trivedi et al. [11] proposed DyRep, a framework that update the representation of a node as it appear in an event in the C-TDG. DyRep captures the continuous-time dynamics leveraging a temporal point process approach. A temporal point process is characterized by the conditional intensity function that models the likelihood of an event to happen given the previous events. DyRep’s conditional intensity function, computed for an event between nodes $u$ and $v$ at time $t$, is:
$\small\lambda_{uv}^{k}(t)=f_{k}(g^{k}_{uv}(t^{-}))$ (20)
@@
$\small f_{k}(z)=\psi_{k}\log\left(1+\exp\left(\frac{z}{\psi_{k}}\right)\right)$ (21)
@@
$\small g^{k}_{uv}(t^{-})=\omega_{k}^{T}\cdot[\mathbf{x}_{u}(t^{-})||\mathbf{x}_{v}(t^{-})]$ (22)
@@
$\small\mathbf{h}_{u}(t)=\sigma(\mathbf{\Theta}_{i}\mathbf{h}_{u}^{loc}(t^{-})+\mathbf{\Theta}_{r}\mathbf{h}_{u}(t^{-}_{u})+\mathbf{\Theta}_{e}(t-t^{-}_{u}))$ (23)
where $h_{u}^{loc}(t^{-})\in\mathbb{R}^{d_{h}}$ is the representation of the aggregation of $u$’s direct neighbors, $t^{-}_{u}$ is the timestamp of the previous event involving node $u$, and $\mathbf{\Theta}_{i},\mathbf{\Theta}_{r}\in\mathbb{R}^{d_{h}\times d_{h}}$ and $\mathbf{\Theta}_{e}\in\mathbb{R}^{d_{h}}$ are learnable parameters. In Equation 23 the first addend propagates neighborhood information, the second self-information, while the third considers the exogenous force that may smoothly update node features during the interval time. To learn $\mathbf{h}_{u}^{loc}(t^{-})$, DyRep uses an attention mechanism similar to the one proposed in the GAT model by Veličković et al. [20]. In this case, the attention coefficient is parametrized by $\mathcal{S}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$, which is a stochastic matrix denoting the likelihood of communication between each pair of nodes. $\mathcal{S}$ is updated according to the conditional intensity function. The aggregated neighborhood representation is
$\small\mathbf{h}_{u}^{loc}(t^{-})=\max(\{\sigma(\alpha_{uv}(t)\cdot\mathbf{h}_{v}(t^{-}))\,|\,v\in\mathcal{N}_{u}^{t})\}),$ (24)
@@ V-2 Stacked architectures
$\small\mathbf{h}^{\ell}_{u}(t)=\text{{mlp}}^{\ell}_{2}(\mathrm{ReLU}(\text{{mlp}}^{\ell}_{1}([\hat{\mathbf{h}}(t)||\mathbf{x}_{u}])))$ (25)
where $\hat{\mathbf{h}}(t)$ is the attentive hidden neighborhood representation obtained as
$\small\begin{split}\mathbf{q}(t)&=\left[\mathbf{Z}(t)\right]_{0}\mathbf{\Theta}_{q}\\ \mathbf{K}(t)&=\left[\mathbf{Z}(t)\right]_{1:n}\mathbf{\Theta}_{K}\\ \mathbf{V}(t)&=\left[\mathbf{Z}(t)\right]_{1:n}\mathbf{\Theta}_{V}\\ \hat{\mathbf{h}}(t)&=\mathrm{attn}(\mathbf{q}(t),\mathbf{K}(t),\mathbf{V}(t))\end{split}$ (26)
where $\mathbf{Z}(t)=\left[\mathbf{x}^{\ell-1}_{u}(t)||\Phi_{d}(0),...,\mathbf{x}^{\ell-1}_{v}(t)||\Phi_{d}(t-t_{v})\right]\in\mathbb{R}^{(d_{h}+d_{t})\times n}$ is the temporal feature matrix that concatenates the representation of each node in the neighborhood of $u$ with the time difference between the current time and the time of the previous event involving the neighbor, with $v\in\mathcal{N}_{u}^{t}$ and $n$ the size of $u$’s neighborhood; $\mathbf{q}(t)$, $\mathbf{K}(t)$, and $\mathbf{V}(t)$ are the query, key and value projections of the matrix; and $\mathrm{attn}$ is an attention mechanism similar to GAT. The dimensional functional mapping $\Phi_{d}:t\rightarrow\mathbb{R}^{d_{t}}$ is defined as
$\small\Phi_{d}(t)=[\cos(\omega_{1}t)\sin(\omega_{1}t),...,\cos(\omega_{d}t)\sin(\omega_{d}t)]$ (27)
@@
$\small\begin{split}m_{u}(t)&=\mathrm{msg}_{src}(\mathbf{s}_{u}(t^{-}),\mathbf{s}_{v}(t^{-}),\Delta t,\mathbf{e}_{uv}(t))\\ m_{v}(t)&=\mathrm{msg}_{dst}(\mathbf{s}_{v}(t^{-}),\mathbf{s}_{u}(t^{-}),\Delta t,\mathbf{e}_{uv}(t)),\end{split}$ (28)
@@
$\small\mathbf{s}_{u}(t)=\mathrm{mem}(\bar{m}_{u}(t),\mathbf{s}_{u}(t^{-}))$ (29)
where $\bar{m}_{u}(t)$ represents the aggregation of computed messages in a batch related to node $u$, and $\mathrm{mem}$ is an RNN. Lastly, the embedding module generates the representation for a node $u$ at time $t$ by exploiting the information stored in the memory module of the node itself and its neighborhood up to time $t$
$\small\mathbf{h}_{u}(t)=\sum_{v\in\mathcal{N}_{u}^{t}}f(\mathbf{s}_{u}(t),\mathbf{s}_{v}(t),\mathbf{x}_{u}(t),\mathbf{x}_{v}(t),\mathbf{e}_{uv})$ (30)
@@ V-3 Random walk based architectures
$\small\mathbb{P}(v)=\frac{\exp(\mathcal{T}(v)-\mathcal{T}(u))}{\sum_{v^{\prime}\in\mathcal{N}^{t}_{u}}\exp(\mathcal{T}(v^{\prime})-\mathcal{T}(u))},$ (31)
@@
$\small\mathbb{P}(v)=\frac{\delta(v,\mathcal{T}(v))}{\sum_{v^{\prime}\in\mathcal{N}^{t}_{u}}\delta(v^{\prime},\mathcal{T}(v^{\prime}))},$ (32)
where $\delta:\mathcal{V}\times\mathbb{R}^{+}\rightarrow\mathbb{Z}^{+}$ is a function that sorts temporal neighbors in descending order depending on time, thus returns a score that biases the distribution towards the selection of edges that are closer in time to the current node.
@@
NeurTWs [67] extends Wang et al. [66] by employing a different sampling strategy for the temporal random walks, which integrates temporal constraints, topological properties, and tree traversals, allowing to sample spatiotemporal-biased random walks. These walks prioritize neighbors with higher connectivity (promoting exploration of more diverse and potentially expressive motifs), while being aware of the importance of recent neighbors. Furthermore, NeurTWs replaces the RNN-based encoding approach for walks with a component based on neural ODEs to facilitate the explicit embedding of irregularly-sampled events.
@@ V-4 Hybrid architectures
$\small\begin{split}\hat{\mathbf{h}}^{\ell}_{u}(t)&=\sum_{v\in\mathcal{N}_{u}^{t}}\text{{mlp}}^{\ell}_{agg}(\mathbf{h}^{\ell-1}_{v}(t)||\mathbf{e}_{uv})\alpha^{-\beta(t-t^{-})}\\ \mathbf{h}^{\ell}_{u}(t)&=\text{{mlp}}^{\ell}_{upd}(\mathbf{h}^{\ell-1}_{u}(t)||\hat{\mathbf{h}}^{\ell}_{u}(t))\end{split}$ (34)
@@ VI The benchmark problem
With this in mind, we designed three benchmarks to assess models that deal with spatio-temopral graphs, general D-TDGs, and C-TDGs. Specifically, we evaluated methods for D-TDGs on both link and node prediction tasks while we constrained the evaluation of C-TDG methods to link prediction tasks due to the scarcity of suitable datasets. To do so, we extended the library PyDGN [69] to the D-TDG learning setting to foster reproducibility and robustness of results. With the same aim, we developed a Pytorch Geometric [70] based framework to allow reproducible results in the continuous scenario. Lastly, in the Supplementary materials (Appendix A) we provide the community with a selection of datasets useful for benchmarking future works. An interest reader is referred to SNAP [71], TSL [72], and Network Repository [73] for a broader data collections.
VI-A Spatio-temopral graph benchmark
In the spatio-temporal setting, we consider three graph datasets for traffic forecasting, i.e., Metr-LA [39], Montevideo [74], and PeMSBay [39], and Traffic [46]. Specifically,
  • Traffic: consists of traffic data collected over a period of three months, with hourly granularity, from 4,438 sensor stations located in the 7th District of California.
We designed each model as a combination of three main components. The first is the encoder which maps the node input features into a latent hidden space; the second is the DGN which computes the spatio-temporal convolution; and the third is a readout that maps the output of the convolution into the output space. The encoder and the readout are MLPs that share the same architecture among all models in the experiments. We performed hyper-parameter tuning via grid search, optimizing the Mean Absolute Error (MAE). We perform a time-based split of the dataset which reserves the first 70% of the data as training set, 15% of the following data as validation set, and the last 15% as test set. We trained the models using Adam optimizer for a maximum of 1000 epochs and early stopping with patience of 50 epochs on the validation error. For each model configuration, we performed 5 training runs with different weight initialization and report the average of the results. We report in Table I the grid of hyper-parameters exploited for this experiment.
TABLE I: The grid of hyper-parameters employed during model selection for the spatio-temporal tasks.
Configs
learning rate $10^{-2}$, $10^{-3}$, $10^{-4}$
weight decay $10^{-3}$, $10^{-4}$
embedding dim 1, 2, 4, 8
$\sigma$ ReLU
Chebishev poly. filter size 1, 2, 3
Results
In Table II we report the results on the spatio-temporal-based experiments.
TABLE II: Mean test scores of the spatio-temporal models and std averaged over 5 random weight initializations. MAE is the optimized metric. The lower, the better.
Montevideo Metr-LA PeMSBay Traffic
Model MAE MSE MAE MSE MAE MSE MAE MSE
A3TGCN 0.3962±${}_{\pm 0.0021}$ 1.0416±${}_{\pm 0.0047}$ 0.3401±${}_{\pm 0.0008}$ 0.3893±${}_{\pm 0.0039}$ 0.2203±${}_{\pm 0.0105}$ 0.2540±${}_{\pm 0.0039}$ 0.1373±${}_{\pm 0.0300}$ 0.0722±${}_{\pm 0.0053}$
DRCNN 0.3499±${}_{\pm 0.0006}$ 1.0686±${}_{\pm 0.0012}$ 0.1218±${}_{\pm 0.0013}$ 0.0960±${}_{\pm 0.0017}$ 0.0569±${}_{\pm 0.0004}$ 0.0398±${}_{\pm 0.0002}$ 0.0153±${}_{\pm 0.0002}$ 0.0018±${}_{\pm 4\cdot 10^{-5}}$
GCRN-GRU 0.3481±${}_{\pm 0.0008}$ 1.0534±${}_{\pm 0.0062}$ 0.1219±${}_{\pm 0.0007}$ 0.0973±${}_{\pm 0.0009}$ 0.0571±${}_{\pm 0.0005}$ 0.0404±${}_{\pm 0.0007}$ 0.0153±${}_{\pm 0.0003}$ 0.0018±${}_{\pm 0.0001}$
GCRN-LSTM 0.3486±${}_{\pm 0.0026}$ 1.0451±${}_{\pm 0.0099}$ 0.1235±${}_{\pm 0.0009}$ 0.0985±${}_{\pm 0.0004}$ 0.0593±${}_{\pm 0.0004}$ 0.0436±${}_{\pm 0.0001}$ 0.0153±${}_{\pm 0.0006}$ 0.0018±${}_{\pm 0.0001}$
TGCN 0.4024±${}_{\pm 0.0022}$ 1.0678±${}_{\pm 0.0049}$ 0.3422±${}_{\pm 0.0046}$ 0.3891±${}_{\pm 0.0058}$ 0.2109±${}_{\pm 0.0039}$ 0.2466±${}_{\pm 0.0033}$ 0.1375±${}_{\pm 0.0299}$ 0.0724±${}_{\pm 0.0052}$
Overall, DCRNN and GCRN-GRU achieve the better performance on the selected tasks. Interestingly, they both rely on Chebyshev spectral convolution and GRU, but with different architectural structure. Indeed, DCRNN employs a stacked architecture, while GCRN-GRU embeds the DGN into the RNN, enabling a combined modeling of the temporal and spatial information. This result shows that there is not a superior architectural design, in these tasks. However, it seems relevant to include a bigger neighborhood in the computation (e.g., by exploiting a larger Chebishev polynomial filter size). Indeed, even if A3TGCN employs an attention mechanism to capture more global information, it is not enough to achieve comparable performance to DCRNN or GCRN-based approaches. Nevertheless, it is noteworthy that the superior performance of these approaches comes at the expense of computational speed, as it is shown in the Supplementary materials (Appendix C).
VI-B D-TDG benchmark
  • Bitcoin-$\alpha$ [78, 79]: a who-trusts-whom network of bitcoin users trading on the platform http://www.bitcoin-alpha.com. To convert this graph into a succession of snapshots, we adopted the same daily aggregation strategy as in [57].
We performed hyper-parameter tuning via grid search, optimizing the Mean Absolute Error (MAE) in the case of node regression, Area Under the ROC curve (AUROC) in the case of link prediction, and balanced accuracy for node classification. We considered the same experimental setting, split strategy, and architectural choice as for the spatio-temporal graphs. In the case of link prediction, we perform negative sampling by randomly sampling non-occurring links from the next future snapshots. We note that in the case of DynGESN, the model employs fixed and randomized weights and only the final readout is trained. We report in Table III the grid of hyper-parameters exploited for this experiment.
TABLE III: The grid of hyper-parameters employed during model selection for the D-TDG tasks. The “***” value refer only to LRGCN model, while $\diamond$” to DynGESN.
Configs
learning rate $10^{-2}$, $10^{-3}$, $10^{-4}$
weight decay $10^{-3}$, $10^{-4}$
embedding dim 8, 16, 32
$\sigma$ ReLU
Chebishev poly. filter size 1, 2, 3
n. bases in the basis-decomposition regularization $*{}^{*}$ None, 10, 30
leakage (i.e., $\gamma$${}^{\diamond}$ 0.1, 0.5, 0.9
random weight init. ${}^{\diamond}$ 0.1, 0.5, 0.9
Results
Table IV shows the results on general D-TDGs. Differently than the spatio-temporal setting, different tasks benefit from different architectures. Indeed, integrating topology’s changes (such as in GCLSTM and LRGCN) is more effective in link prediction tasks, while evolving the parameters of the DGN is more beneficial for node-level tasks, since it is more difficult to change the parameters of a static DGN to predict the topological evolution of the system. Notably, DynGESN achieves comparable results by exploiting only few trainable parameters and reduced computational overhead (see Supplementary materials, Appendix C), showing an advantageous tradeoff between performance and complexity. This makes it an ideal choice when the computational resources are limited.
TABLE IV: Mean test scores and std of DGNs for general D-TDGs averaged over 5 random weight initializations. For MAE and MSE scores, lower values corresponds to better performances, while for Balacend Acc, AUROC and F1 the higher values are better. In red is reported the optimized metric.
Node-level tasks Link-level tasks
Twitter tennis Elliptic AS-773 Bitcoin $\alpha$
MAE MSE Balanced Acc F1 AUROC F1 AUROC F1
DynGESN 0.1944±${}_{\pm 0.0056}$ 0.3708±${}_{\pm 0.0411}$ 50.56±${}_{\pm 1.10}$ 79.2±${}_{\pm 19.62}$ 95.34±${}_{\pm 0.04}$ 79.83±${}_{\pm 5.27}$ 97.68±${}_{\pm 0.12}$ 69.98±${}_{\pm 1.57}$
EvolveGCN-H 0.1735±${}_{\pm 0.0007}$ 0.2858±${}_{\pm 0.0074}$ 49.52±${}_{\pm 1.55}$ 92.54±${}_{\pm 8.39}$ 59.52±${}_{\pm 17.53}$ 39.85±${}_{\pm 34.24}$ 51.35±${}_{\pm 2.88}$ 29.55±${}_{\pm 30.58}$
EvolveGCN-O 0.1749±${}_{\pm 0.0007}$ 0.3020±${}_{\pm 0.0111}$ 49.23±${}_{\pm 1.03}$ 90.80±${}_{\pm 12.67}$ 58.90±${}_{\pm 17.80}$ 29.99±${}_{\pm 37.10}$ 51.42±${}_{\pm 2.84}$ 31.74±${}_{\pm 29.98}$
GCLSTM 0.1686±${}_{\pm 0.0015}$ 0.2588±${}_{\pm 0.0049}$ 48.20±${}_{\pm 1.80}$ 70.84±${}_{\pm 30.01}$ 96.35±${}_{\pm 0.01}$ 91.22±${}_{\pm 0.13}$ 97.75±${}_{\pm 0.17}$ 91.22±${}_{\pm 1.38}$
LRGCN 0.1693±${}_{\pm 0.0014}$ 0.2507±${}_{\pm 0.0057}$ 47.84±${}_{\pm 3.37}$ 65.69±${}_{\pm 20.21}$ 94.77±${}_{\pm 0.23}$ 89.59±${}_{\pm 0.33}$ 98.05±${}_{\pm 0.03}$ 91.33±${}_{\pm 0.08}$
VI-C C-TDG benchmark
We report in Table V the grid of hyper-parameters exploited for this experiment.
TABLE V: The grid of hyper-parameters employed during model selection for the C-TDG tasks.
Configs
learning rate $10^{-3}$,$10^{-4}$
weight decay $10^{-4}$,$10^{-5}$
n. DGN layers 1, 3
embedding dim 32, 64, 96
DGN dim emb dim, emb dim / 2
$\sigma$ tanh
Neighborhood sampler size 5
Results
We report the results of the C-TDG experiments in Table VI.
TABLE VI: Mean test scores and std of DGNs for C-TDGs averaged over 5 random weight initializations. The higher, the better. The models are trained to maximize the AUROC score.
Wikipedia Reddit LastFM
Model AUROC F1 AUROC F1 AUROC F1
EdgeBank 91.82 91.09 96.42 96.29 94.72 94.43
Overall, TGN generally outperforms all the other methods, showing consistent improvements over DyRep and JODIE. This result shows how the spatial information is fundamental for the effective resolution of the tasks. Indeed, an advantage of TGAT and TGN is that they can exploit bigger neighborhoods with respect to DyRep, which uses the information coming from one-hop distance, and JODIE, which only encode the source and destination information. Despite these results, we observe that the temporal information is still extremely relevant to achieve good performance. In fact, the EdgeBank baseline is able to exceed 91% AUROC score by only looking at the graph’s history. This is even more evident in the LastFM task, which, as observed in [80], contains more reoccurring edges with respect to Wikipedia and Reddit. Consequently, such a task is comparatively easier to solve by solely exploiting these temporal patterns. Considering that EdgeBank’s performance is directly correlated to the number of memorized edges, in this task, it is able to outperform all the other methods. Lastly, it is worth mentioning that the enhanced performance of TGN and TGAT is accompanied by a trade-off in computational speed, as illustrated in the Supplementary materials (Appendix C).
VII Conclusions
In order to further improve the maturity of representation learning for dynamic graphs, we believe that certain aspects should be deepened. A future interesting direction, in this sense, is to extend the work that has been done for heterophilic (static) graphs [81, 82, 83] to the temporal domain. This will require addressing the problem of generating information-rich node representations when neighboring nodes tend to belong to different classes. A similar challenge is the one of heterogeneus graphs [84, 85], which contain different types of nodes and links. In this scenario, new architectures should learn the semantic-level information coming from node and edge types, in addition to topological and label information. While these are interesting future directions, we observe that there are compelling challenges that need addressing and that relate to studying, in the temporal domain, aspects such as robustness to adversarial attacks [86, 87], over-smoothing [88] (a phenomenon where all node features become almost indistinguishable after few embedding updates), over-squashing [89, 90] (phenomenon preventing DGNs to propagate and preserve long-term dependencies between nodes), TDGs’ expressive power [23, 91], and whole graph learning [92].
Acknowledgements
Supplementary Material of
“Deep learning for dynamic graphs: models and benchmarks”
-A Datasets, Models, and previous studies
In Table VII we provide the community with a selection of datasets useful for benchmarking future works. In Table VIII we report an overview of the examined models with respect to the specific changes in the graph structure that each model was designed to address, i.e., node/edge additions/deletions. We observe that each method is designed to address changes in node/edge features. Furthermore, the methods developed for D-TDGs that are not specifically designed to address changes in the node set can still be applied to tasks involving an evolving node set by treating nodes not in the current snapshot as isolated entities. Lastly, in Table IX we show a comparative analysis with previous benchmarking studies and surveys, assessing the provision of datasets and benchmarks and delineating the types of analyzed graphs.